OpenStudy (daryan):

A store sells tents, sleeping bags, and camp stools. A customer buys a tent, 5 sleeping bags and 3 camp stools for $900. The price of the tent is 7 times the cost of a camp stool. The cost of a sleeping bag is $30 more than the cost of a camp stool. Find the cost of the camp stool.

OpenStudy (anonymous):

Let t, s and c be the cost of a tent, sleeping bag and a camp stool respectively. Solve the following equations which are the result of translating the problem text to math equations using the definitions in the first sentence. \[\{3 c+5 s+t=900,t=7 c,s=c+30\} \]\[\{c=50,s=80,t=350\} \]